# Parameter-Efficient Continual Fine-Tuning: A Survey

### v1

@@ 4.3 Selective Methods
The family of selective methods for PEFT includes such techniques that aim to pick and finetune only a small subset of the model parameters, hence reducing computational requirements and memory footprint.
Masking [159] starts from the idea of selecting the weights that are important for the downstream task, instead of finetuning the entire model. Based on the lottery ticket hypothesis [33], they propose to select the parameters with a series of learned binary masks, one for each task. The authors experimented with Transformer-based architectures, such as BERT [23], RoBERTa [88] and DistilBERT [120], on different NLP tasks, including part-of-speech tagging, named-entity recognition, sequence classification, and reading comprehension. In this approach, for each weight matrix $\textbf{W}^{l}$ of the $l$-th transformer block, they randomly initialize a matrix $\textbf{M}^{l}$, which is then binarized via an element-wise thresholding function. The binary mask is then applied to the weights by Hadamard product:
@@
BitFit [154] is a sparse finetuning method, in which the basic concept is to update only the bias terms, while keeping the rest of the network frozen. Such approach is very parameter-efficient, since very few weights are modified: for BERT architecture, they amount to less than 0.1% of the total number of the parameters. In particular, the authors consider the bias terms of key, query and value weights of each self-attention head, for each layer of the BERT encoder, plus the bias parameters from the MLP layers on top.
The BitFit approach was evaluated against the GLUE benchmark [136], obtaining comparable results to naive finetuning, but modifying a tiny portion of the weights. Moreover, they experimented with updating a subset of the bias parameters, achieving only slightly worse results.
@@ 5 The Confluence of CL and PEFT
5.1 PECFT Methodologies
In recent times, several works in CL [143, 144, 125, 113] have incorporated different PEFT methods, each offering a different complementary advantage over its CL counterpart. In this survey, we split these PECFT method based on the kind of PEFT method applied.
5.1.1 Adapter-based PECFT
@@
Similarly, Adapter-based Continual Learning (ACL) [157] introduces a framework that uses a feature adapter to combat catastrophic forgetting. It is designed with a task-specific head, which cleverly groups all previously learned classes into a single “out-of-distribution” category, enabling more effective feature discrimination. The approach also includes a collaborative fine-tuning mechanism, ensuring that the outputs of the different classifiers remain comparable, facilitating accurate head selection during inference. The framework’s effectiveness is validated through comprehensive testing on three image datasets, demonstrating ACL’s superior performance in continually learning new classes while preserving existing knowledge.
In adapter-based CL frameworks, routing mechanisms enable dynamic selection of specialized adapters, facilitating efficient knowledge accumulation and modular model expansion. [151] introduces a framework that dynamically expands pre-trained CLIP models using Mixture-of-Experts (MoE) adapters, featuring a Distribution Discriminative Auto-Selector that intelligently routes inputs between the MoE Adapter and original CLIP model. This architecture demonstrates remarkable efficiency, reducing parameter training overhead by 60% while maintaining competitive performance.
The Continual Adapter Tuning (CAT) [15] tackles Aspect Sentiment Classification (ASC) by combining task-specific adapters with a frozen pre-trained backbone. CAT uses its continual adapter initialization technique for knowledge transfer between tasks and label-aware contrastive learning to jointly optimize features and classifiers. The framework eliminates the need for task IDs during inference by employing a majority voting strategy across adapter paths. Through validation on 19 ASC datasets, CAT demonstrates state-of-the-art performance in maintaining effectiveness across domains while enabling efficient knowledge transfer.
ATLAS [76] proposes a two-stage learning paradigm that effectively balances the preservation of existing knowledge with the acquisition of new capabilities by incorporating vector learning method, which intelligently combines information from different adapters based on cosine similarity between tasks. Unlike previous parameter-efficient solutions that created isolated modules for each task, ATLAS minimizes knowledge redundancy while expanding the model’s representational capacity. It employes both multi-modal and uni-modal tasks in upstream continual learning, providing valuable insights into how multi-modal model updates affect performance across different modalities. Its ability to enhance distribution richness and improve generalization capability makes it a promising solution for CL in Vision-Language models.
The Expand and Merge [51] framework introduces a parameter-efficient architecture that combines adapter layers with Vision-Language models. The approach is built on two key innovations: first, it employs specially designed adapter layers that expand for new tasks while keeping old knowledge intact through frozen parameters; second, it leverages a pretrained text encoder’s fixed embedding space to guide the vision encoder’s continual learning process through vision-language pretraining models like CLIP. To effectively manage knowledge integration, the framework implements an adaptive fusion mechanism using scaling weights at different network depths, complemented by a unique parameter merging stage that prevents performance degradation while controlling parameter growth. This design addresses two common limitations in the field: the parameter bloat, typical of expansion-based approaches, and the over-constraining of new learning found in regularization-based methods. Through extensive validation across three datasets, the framework demonstrates superior performance compared to current state-of-the-art methods, exhibiting robust performance maintenance on both old and new tasks while effectively managing parameter efficiency.
5.1.2 LoRA-Based PECFT
@@
InfLoRA [79] is a reparameterization technique that confines weight updates to a carefully designed subspace, eliminating interference between new and old tasks. This subspace-constrained learning improves upon existing approaches that either reuse parameters or randomly expand them for new tasks. The method approximates the gradient space using the input matrix of new tasks, making it particularly effective in CIL scenarios where task IDs are unavailable during inference. Through this carefully designed subspace and branch expansion architecture, InfLoRA demonstrates superior performance compared to state-of-the-art methods, particularly in maintaining stability on old tasks while effectively adapting to new ones.
@@
Prototype Guided Incremental LoRA (PILoRA) [37] addresses two key challenges in federated class incremental learning: catastrophic forgetting and data heterogeneity across clients. The method combines prototype learning with parameter-efficient fine-tuning through an innovative two-pronged approach. First, it introduces incremental LoRA, which mitigates forgetting by constraining different learning stages to orthogonal subspaces, allowing efficient knowledge accumulation through parameter summation during inference. Second, it implements a prototype re-weight module that leverages heuristic information between prototypes and class features to address classifier bias without retraining. Built on the ViT backbone, PILoRA performs well on standard benchmarks while maintaining strong performance even under extreme heterogeneous conditions where other methods falter. Its effectiveness stems from the unified approach to both continual learning and federated learning challenges, recognizing their shared need for robust feature representations and bias mitigation in classifiers. While InfLoRA focuses on confining weight updates to carefully designed subspaces to prevent interference between tasks, PILoRA expands on this idea by implementing orthogonal subspaces for different learning stages while adding prototype learning to handle data heterogeneity.
@@
A slightly different approach comes from Dual Low-Rank Adaptation (DualLoRA) [14]. It employs two distinct low-rank adapters working in concert: an orthogonal adapter that operates in subspaces perpendicular to previous task features to preserve old knowledge, and a residual adapter that learns in task-specific subspaces to enable effective adaptation to new tasks. DualLoRA is quite efficient in feature subspace extraction; it uses Singular Value Decomposition, which eliminates the need for multiple forward passes that burden previous approaches like InfLoRA. During inference, DualLoRA uses its dynamic memory mechanism which modulates the residual adapter’s output based on task relevance computed from input samples. This dynamic adjustment not only improves feature embeddings but also enables accurate task identification without explicit task labels. By maintaining parameter efficiency while offering superior performance across multiple benchmarks, DualLoRA represents a significant advance over existing PEFT methods for CL.
Experience Replay Informative-Low Rank Adaptation (ERI-LoRA) [7] tackles CL in task-oriented dialogue systems by combining replay-based methods with parameter-efficient fine-tuning. The method improves upon standard experience replay by introducing a weighted distribution sampling strategy for previous domain examples, while leveraging LoRA for efficient parameter updates on LLaMA2. Instead of random sample selection, ERI-LoRA employs an informed sampling approach that considers class label distributions. Through extensive evaluation on intent detection and slot-filling tasks across multiple datasets and orderings, ERI-LoRA shows significant improvements over state-of-the-art methods, achieving a 13% higher F1 score in slot-filling and 0.85% better accuracy in intent detection.
Another replay-based method is Task Arithmetic with LoRA for CL [19]. Instead of sequential training, that brings forgetting risks, this approach trains individual LoRA modules for each task separately, then it combines the task vectors using task arithmetic rules before merging them back into the pre-trained Vision Transformer. The method’s innovation lies in its efficient use of resources by only training small LoRA modules and leveraging a minimal memory buffer of 10 samples per class for final fine-tuning.
CL with STack-And-Mask INcremental Adapters (STAMINA) [124] enhances text-to-image diffusion models’ ability to learn long sequences of concepts without forgetting. The method combines LoRA with two key technical innovations: hard-attention masks parameterized by low-rank MLPs using Gumbel softmax, and learnable MLP tokens that replace traditional custom token embeddings, with all trainable parameters capable of being folded back into the base model after training, eliminating inference overhead. It is also able to maintain plasticity thanks to its sparse adaptation approach. Through comprehensive evaluation on a 50-concept benchmark of landmarks and faces, STAMINA demonstrates superior performance over existing methods while requiring fewer training steps to achieve concept mastery.
@@
Interpolation-based LoRA [115] approaches CL in LLMs through mode connectivity, i.e. the observation that different optimization minima can be connected through low-loss valleys. Rather than viewing catastrophic forgetting as a binary trade-off, it implements a dual-memory architecture: a fast learner that rapidly adapts to new tasks — which provides plasticity — and a slow learner that consolidates long-term knowledge — maintaining stability. This design is inspired by empirical findings that mode connectivity exists in parameter-efficient fine-tuning of LLMs, and that linear interpolation between task-specific optima can achieve better plasticity-stability trade-offs than previous approaches. [115] tries to solve continual learning in a rather different way: instead of trying to preserve discrete task knowledge, it leverages the continuous nature of the parameter space to find optimal interpolation points that serve both past and present learning objectives.
5.1.3 Prompt-based PECFT
@@
Learning to Prompt (L2P) [144] is one of the first works that applies prompting to CL. In this work, a prompt is trained for each task on a pre-trained network and subsequently stored in a key-value prompt pool, where the key is learnable. To select the most adequate prompts at inference time for a given input, the authors introduce a deterministic query function, i.e. the pre-trained network, to extract features from the input. Prompts are then selected based on the similarity between input query and prompt keys. L2P was evaluated against the main CL strategies, showing enhanced performance in class-incremental, domain-incremental and task-agnostic scenarios. Similarly, DualPrompt [143] builds on the prompt pool concept by introducing two types of prompts: general prompts — for common features across tasks — and expert prompts — for task specific instructions. These prompts are attached to two distinct groups of contiguous attention layers, selected through heuristic search. DualPrompt leverages the L2P workflow: each e-prompt is associated with a task-specific key that is learned to match the input features; at inference time, a query function on the test sample is used to retrieve the best e-prompt. Compared against leading memory-based and architecture-based CL methods, DualPrompts outperforms all while remaining rehearsal-free. It also beats L2P by a modest margin.
An additional technique for prompt-based approaches is called Language Guidance for Prompt-based Continual Learning (LGCL) [64]. In this method, the underlying assumption is that natural language can be used as a shared representation space for tasks with different visual features. Hence, the authors propose to employ a pre-trained text encoder to obtain a representation of both tasks and single classes. Then, the task-level representation is used to select the prompts via cosine similarity with each key, while the class-level embedding is encoded in the output feature of the model, to perform the classification. LGCL is shown to consistently outperform CL baselines on both Split-CIFAR100 [69] and Split-ImageNet-R [143].
Another study that builds on the prompt pool concept is the Prompt Gradient Projection (PGP) approach [109]. It represents the first work that analyzes anti-forgetting mechanisms integrated with prompt tuning techniques. In particular, the authors employ the Gradient Projection Method [118], which demonstrates theoretically that forgetting is mitigated if the weights are updated in the orthogonal direction to the subspace spanned by the previous features. PGP was evaluated in combination with L2P and DualPrompt, showing improving results of accuracy and forgetting.
@@
A slight variation from L2P comes from S-Prompts [140]. This approach is specifically applied to Domain-Incremental Learning scenarios, without the need to keep a buffer of samples from previous tasks. In L2P and similar works the prompts are shared among tasks, meaning that the new knowledge shares the same feature space of the old tasks, limiting the learning capacity and possibly inducing interference. Conversely, S-Prompts instantiate a new prompt for each domain, so that each prompt can obtain optimal performance on its task independently from the other prompts. At inference time, K-NN is employed to find the closest domain to a given test sample, and the corresponding prompt is prepended to the input to perform classification. In the paper, the authors show an exhaustive comparison with both prompting methods and standard CL baselines, showing great advancements on the state-of-the-art results.
@@
A limitation of the methods we have discussed so far consists in the fixed size both of the prompt pool and the prompts themselves. This limits their capability to scale, being difficult to increase the learning capacity when needed. CODA-Prompt [125] overcomes this issue defining a set of prompt components: for each input sample, the components are merged via weighted summation, forming a so-called decomposed prompt. By doing so, the method can automatically adjust the prompting capacity depending on the complexity of the task at hand. Furthermore, the authors introduced an attention mechanism on the input query, which allows to focus on relevant features only, reducing also the input dimensionality. Unlike L2P and DualPrompt, CODA-Prompt is optimized in an end-to-end fashion, using directly the classification loss, which helps increasing performance and establishing new SOTA results. Indeed, this method surpasses DualPrompt on both class-incremental and domain-incremental scenarios, especially when the number of tasks increases.
@@
A different approach comes from Progressive Prompts [113]. In this work, a new prompt is learned for each new task, and it is then concatenated with the previously learned ones. Each prompt $P_{k}$ is trained only during corresponding task $T_{k}$, while it is kept frozen during subsequent tasks. The prompts concatenation is then prepended to the input embeddings. This approach was evaluated on two CL benchmarks for text classification using two widely-used language models. The results demonstrate its effectiveness in improving performance and mitigating catastrophic forgetting, by facilitating greater forward transfer across tasks.
@@
Following the idea of prompts concatenation, Prompt of Prompts (POP) [50], proposes an approach involving two sets of prompts. First, one or more prompts are learned per each task, with the objective to discriminate among the classes in the task. Such prompts are then kept frozen for subsequent tasks, but are still given in input to the model; in this way, the new prompts are forced to learn only the new information from the current task, without duplicating knowledge present in other prompts. Additionally, this method uses a second set of prompts, called Prompt of Prompts, to combine the representations from different tasks. While the task prompts are frozen after training on the corresponding task, the Prompt of Prompts set is learned continually, so to integrate and update the features across all the tasks.

### v2

@@ 4.3 Selective Methods
The family of selective methods for PEFT includes such techniques that aim to pick and finetune only a small subset of the model parameters, hence reducing computational requirements and memory footprint. Masking [170] starts from the idea of selecting the weights that are important for the downstream task, instead of finetuning the entire model. Based on the lottery ticket hypothesis [34], they propose to select the parameters with a series of learned binary masks, one for each task. The authors experimented with Transformer-based architectures, such as BERT [24], RoBERTa [94] and DistilBERT [128], on different NLP tasks, including part-of-speech tagging, named-entity recognition, sequence classification, and reading comprehension. In this approach, for each weight matrix $\textbf{W}^{l}$ of the $l$-th transformer block, they randomly initialize a matrix $\textbf{M}^{l}$, which is then binarized via an element-wise thresholding function. The binary mask is then applied to the weights by Hadamard product:
@@
BitFit [165] is a sparse finetuning method, in which the basic concept is to update only the bias terms, while keeping the rest of the network frozen. Such approach is very parameter-efficient, since very few weights are modified: for BERT architecture, they amount to less than 0.1% of the total number of the parameters. In particular, the authors consider the bias terms of key, query and value weights of each self-attention head, for each layer of the BERT encoder, plus the bias parameters from the MLP layers on top. The BitFit approach was evaluated against the GLUE benchmark [145], obtaining comparable results to naive finetuning, but modifying a tiny portion of the weights. Moreover, they experimented with updating a subset of the bias parameters, achieving only slightly worse results.
@@ 5 The Confluence of CL and PEFT
In the following sections, we provide a detailed overview of the main PECFT methodologies, categorizing them according to the type of PEFT strategy they employ. A visual taxonomy of these methods is presented in Figure 2.
Figure 2: Taxonomy of Parameter-Efficient Continual Fine-Tuning methods
5.1 Adapter-based PECFT
5.1.1 Routing and Selection Mechanisms
In adapter-based CL frameworks, routing mechanisms enable dynamic selection of specialized adapters, facilitating efficient knowledge accumulation and modular model expansion. Mixture-of-Experts (MoE) Adapters [162] introduces a framework that dynamically expands pre-trained CLIP models. This approach features a Distribution Discriminative Auto-Selector that intelligently routes inputs between the MoE Adapter and the original CLIP model. This architecture demonstrates remarkable efficiency, reducing parameter training overhead by 60% while maintaining competitive performance.
5.1.2 Merging and Fusion Strategies
ATLAS [78] proposes a two-stage learning paradigm that effectively balances the preservation of existing knowledge with the acquisition of new capabilities by incorporating vector learning method, which intelligently combines information from different adapters based on cosine similarity between tasks. Unlike previous parameter-efficient solutions that created isolated modules for each task, ATLAS minimizes knowledge redundancy while expanding the model’s representational capacity. It employs both multi-modal and uni-modal tasks in upstream continual learning, providing valuable insights into how multi-modal model updates affect performance across different modalities. Its ability to enhance distribution richness and improve generalization capability makes it a promising solution for CL in Vision-Language models.
The Expand and Merge [54] framework introduces a parameter-efficient architecture that combines adapter layers with Vision-Language models. The approach is built on two key innovations: first, it employs specially designed adapter layers that expand for new tasks while keeping old knowledge intact through frozen parameters; second, it leverages a pretrained text encoder’s fixed embedding space to guide the vision encoder’s continual learning process through vision-language pretraining models like CLIP. To effectively manage knowledge integration, the framework implements an adaptive fusion mechanism using scaling weights at different network depths, complemented by a unique parameter merging stage that prevents performance degradation while controlling parameter growth. This design addresses two common limitations in the field: the parameter bloat, typical of expansion-based approaches, and the over-constraining of new learning found in regularization-based methods.
5.1.3 Task Specific Strategies
@@
Similarly, Adapter-based Continual Learning (ACL) [168] introduces a framework that uses a feature adapter to combat catastrophic forgetting. It is designed with a task-specific head, which groups all previously learned classes into a single “out-of-distribution” category, enabling more effective feature discrimination. The approach also includes a collaborative fine-tuning mechanism, ensuring that the outputs of the different classifiers remain comparable, facilitating accurate head selection during inference.
The Continual Adapter Tuning (CAT) [15] tackles Aspect Sentiment Classification (ASC) by combining task-specific adapters with a frozen pre-trained backbone. CAT uses its continual adapter initialization technique for knowledge transfer between tasks and label-aware contrastive learning to jointly optimize features and classifiers. The framework eliminates the need for task IDs during inference by employing a majority voting strategy across adapter paths.
5.1.4 Subspace-based Strategies
Expandable Subspace Ensemble (EASE) [173] addresses task interference by creating task-specific subspaces through lightweight adapter expansion. Unlike methods that modify shared parameters, EASE trains a distinct adapter for each new task, enabling conflict-free learning where new tasks don’t harm previous knowledge. The method concatenates embeddings from all task-specific subspaces to create high-dimensional representations for holistic decision-making. EASE’ utilizes a semantic-guided prototype complement strategy that synthesizes old class prototypes in new subspaces without exemplars. It extracts class-wise similarity in co-occurrence spaces and uses this semantic information to reconstruct missing prototypes. During inference, subspace reweighting highlights contributions from matching task-specific adapters.
5.2 LoRA-Based PECFT
@@ 5.2.1 Orthogonal Subspace Methods
InfLoRA [82] is a reparameterization technique that confines weight updates to a carefully designed subspace, eliminating interference between new and old tasks. This subspace-constrained learning improves upon existing approaches that either reuse parameters or randomly expand them for new tasks. The method approximates the gradient space using the input matrix of new tasks, making it particularly effective in CIL scenarios where task IDs are unavailable during inference.
@@
Prototype Guided Incremental LoRA (PILoRA) [39] addresses two key challenges in federated class incremental learning: catastrophic forgetting and data heterogeneity across clients. The method combines prototype learning with parameter-efficient fine-tuning through an innovative two-pronged approach. First, it introduces incremental LoRA, which mitigates forgetting by constraining different learning stages to orthogonal subspaces, allowing efficient knowledge accumulation through parameter summation during inference. Second, it implements a prototype re-weight module that leverages heuristic information between prototypes and class features to address classifier bias without retraining. Its effectiveness stems from the unified approach to both continual learning and federated learning challenges, recognizing their shared need for robust feature representations and bias mitigation in classifiers. While InfLoRA focuses on confining weight updates to carefully designed subspaces to prevent interference between tasks, PILoRA expands on this idea by implementing orthogonal subspaces for different learning stages while adding prototype learning to handle data heterogeneity.
SD-LoRA[156] introduces a decoupled low-rank adaptation method for class-incremental learning that explicitly separates the learning of direction and magnitude in LoRA updates. After learning the low-rank direction matrices during the first task, SD-LoRA freezes them and optimizes only the task-specific magnitude vectors for subsequent tasks. This design ensures that all tasks share a common representational subspace, reducing interference and eliminating the need for rehearsal or dynamic routing. The approach is fully end-to-end, does not expand the model over time, and supports task-agnostic inference using a single unified adapter.
@@
A slightly different approach comes from Dual Low-Rank Adaptation (DualLoRA) [14]. It employs two distinct low-rank adapters that work in concert: an orthogonal adapter, which operates in subspaces perpendicular to previous task features, to preserve old knowledge, and a residual adapter, which learns in task-specific subspaces to enable effective adaptation to new tasks. DualLoRA is quite efficient in feature subspace extraction, as it utilises Singular Value Decomposition, which eliminates the need for multiple forward passes that burden previous approaches, such as InfLoRA. During inference, DualLoRA uses its dynamic memory mechanism which modulates the residual adapter’s output based on task relevance computed from input samples. This dynamic adjustment not only enhances feature embeddings but also facilitates accurate task identification without requiring explicit task labels.
LoRAC-IPC [85], on the other hand, addresses the challenge of parameter interference in LoRA-based continual learning. The method introduces a two-stage approach: Importance-based Parameter Consolidation (IPC) and Orthogonal LoRA Composition (LoRAC). In the first stage, LoRAC-IPC identifies critical LoRA parameters from earlier tasks—those contributing most to past performance—and freezes them. In the second stage, it ensures that new LoRA updates are orthogonal to frozen directions using QR decomposition, effectively preventing destructive interference. Unlike naive LoRA stacking or merging, this composition maintains compatibility across tasks
5.2.2 Merging and Fusion Strategies
@@
Interpolation-based LoRA [123] approaches CL in LLMs through mode connectivity, i.e. the observation that different optimization minima can be connected through low-loss valleys. Rather than viewing catastrophic forgetting as a binary trade-off, it implements a dual-memory architecture: a fast learner that rapidly adapts to new tasks, which provides plasticity, and a slow learner that consolidates long-term knowledge, maintaining stability. This design is inspired by empirical findings that mode connectivity exists in parameter-efficient fine-tuning of LLMs, and that linear interpolation between task-specific optima can achieve better plasticity-stability trade-offs than previous approaches.
MoLE [155] introduces a hierarchical mixture-of-experts framework for composing multiple LoRA adapters without retraining the base model. Rather than using naive arithmetic merging, MoLE treats each LoRA as an expert and learns layer-wise softmax-normalized gating weights to combine their contributions. Specifically, for each transformer layer, the LoRA updates from all adapters are merged into a single update via a learned weighted sum, where the weights are unique to each layer. This approach allows fine-grained control over how task-specific knowledge is integrated, enabling modular, task-agnostic inference without the need for retraining or access to task IDs.
Merging-based approach is used in Task Arithmetic with LoRA for CL [19]. Instead of sequential training, which introduces forgetting risks, this approach trains individual LoRA modules for each task separately. It then combines the task vectors using task-specific arithmetic rules before merging them back into the pre-trained Vision Transformer. The method’s innovation lies in its efficient use of resources, as it only trains small LoRA modules and leverages a minimal memory buffer of 10 samples per class for final fine-tuning.
5.2.3 Masking-based Strategies
CL with STack-And-Mask INcremental Adapters (STAMINA) [132] enhances text-to-image diffusion models’ ability to learn long sequences of concepts without forgetting. The method combines LoRA with two key technical innovations: hard-attention masks parameterized by low-rank MLPs using Gumbel softmax, and learnable MLP tokens that replace traditional custom token embeddings, with all trainable parameters capable of being folded back into the base model after training, eliminating inference overhead. It is also able to maintain plasticity thanks to its sparse adaptation approach.
5.3 Prompt-based PECFT
5.3.1 Prompt Pool-based Strategies
@@
Learning to Prompt (L2P) [153] is one of the first works that applies prompting to CL. In this work, a prompt is trained for each task on a pre-trained network and subsequently stored in a key-value prompt pool, where the key is learnable. To select the most adequate prompts at inference time for a given input, the authors introduce a deterministic query function, i.e. the pre-trained network, to extract features from the input. Prompts are then selected based on the similarity between the input query and prompt keys. L2P was evaluated against the main CL strategies in class-incremental, domain-incremental, and task-agnostic scenarios.
Similarly, DualPrompt [152] builds on the prompt pool concept by introducing two types of prompts: general prompts, for common features across tasks, and expert prompts, for task-specific instructions. These prompts are attached to two distinct groups of contiguous attention layers, selected through heuristic search. DualPrompt leverages the L2P workflow: each expert prompt is associated with a task-specific key that is learned to match the input features; at inference time, a query function on the test sample is used to retrieve the best e-prompt.
@@
A slight variation from L2P comes from S-Prompts [149]. This approach is specifically applied to Domain-Incremental Learning scenarios, without the need to keep a buffer of samples from previous tasks. In L2P and similar works, the prompts are shared among tasks, meaning that the new knowledge shares the same feature space of the old tasks, limiting the learning capacity and possibly inducing interference. Conversely, S-Prompts instantiate a new prompt for each domain, so that each prompt can obtain optimal performance on its task independently from the other prompts. At inference time, K-NN is employed to find the closest domain to a given test sample, and the corresponding prompt is prepended to the input to perform classification.
Language Guidance for Prompt-based Continual Learning (LGCL) [67] introduces a method for selecting prompts using semantic information derived from natural language. It uses a pre-trained text encoder to generate task- and class-level embeddings. Task embeddings are matched to prompt keys via cosine similarity to guide prompt selection, while class embeddings are injected into the output features for classification.
5.3.2 Prompt Composition and Sequential Learning
@@
One drawback of prompt-based methods we have considered up to this point lies in the fixed dimensions of both the prompt pool and the prompts themselves. This limits their capability to scale, making it difficult to increase the learning capacity when needed. CODA-Prompt [133] overcomes this issue by defining a set of prompt components: for each input sample, the components are merged via weighted summation, forming a so-called decomposed prompt. By doing so, the method can automatically adjust the prompting capacity depending on the complexity of the task at hand. Furthermore, the authors introduced an attention mechanism on the input query, which allows focusing on relevant features only, reducing also the input dimensionality. CODA-Prompt is optimized in an end-to-end fashion, using directly the classification loss, which helps increase performance.
@@
Progressive Prompts [121] offers a more direct approach. In this work, a new prompt is learned for each new task, and it is then concatenated with the previously learned ones. Each prompt $P_{k}$ is trained only during the corresponding task $T_{k}$, while it is kept frozen during subsequent tasks. The prompts’ concatenation is then prepended to the input embeddings. This approach was evaluated on two CL benchmarks for text classification using two widely used language models. The results demonstrate its effectiveness in improving performance and mitigating catastrophic forgetting by facilitating greater forward transfer across tasks.
@@
Following the idea of prompts concatenation, Prompt of Prompts (POP) [53] proposes an approach involving two sets of prompts. First, one or more prompts are learned for each task, with the objective of distinguishing between classes in the task. Such prompts are then kept frozen for subsequent tasks, but are still given in input to the model; in this way, the new prompts are forced to learn only the new information from the current task, without duplicating knowledge present in other prompts. Additionally, this method uses a second set of prompts, called Prompt of Prompts, to combine the representations from different tasks. While the task prompts are frozen after training on the corresponding task, the Prompt of Prompts set is learned continually, so as to integrate and update the features across all the tasks.
5.3.3 Orthogonal Subspace Approaches
Prompt Gradient Projection (PGP) approach [117]. represents the first work that analyzes anti-forgetting mechanisms integrated with prompt tuning techniques. In particular, the authors employ the Gradient Projection Method [126], which demonstrates theoretically that forgetting is mitigated if the weights are updated in the orthogonal direction to the subspace spanned by the previous features.
Visual Prompt Tuning (VPT) [98] builds on prompt-based continual learning by addressing the interference introduced during prompt updates in vision transformers. The authors hypothesize that naively updating prompts across tasks leads to inconsistent gradient directions and degrades performance. To counter this, VPT projects the prompt update into the null space of previously learned prompt gradients, ensuring that new learning has minimal impact on prior knowledge. Unlike earlier works, VPT‑NSP takes into account ViT-specific components such as LayerNorm and self-attention when computing the projection basis, making it better aligned with the model’s internal representations. Experimental results show that enforcing gradient orthogonality at the prompt level is effective in reducing forgetting in ViT-based continual learning.
5.4 Unified Frameworks
Unlike most PEFT-based continual learning methods that focus on a single technique, such as adapters, prompts, or LoRA, some recent works propose unified frameworks that offer flexibility in applying various PEFT strategies within a continual learning setting. These frameworks aim to decouple the continual learning strategy from the specific choice of PEFT module, enabling broader applicability and easier experimentation.
APER (AdaPt and mERge) [171] presents a unified framework that can be orthogonally combined with any parameter-efficient tuning method to address class-incremental learning. The method operates through a two-stage protocol: first adapting the pre-trained model using any PEFT technique (VPT-Deep/Shallow, Scale & Shift, Adapter modules, or full fine-tuning) on only the first incremental task, then concatenating the embeddings from both the adapted model and the original frozen PTM to create unified representations. This concatenated embedding function remains frozen throughout subsequent learning stages, with new classes handled through prototype-based classifiers that leverage both the task-specific adaptivity of the fine-tuned model and the broad generalizability of the original PTM. A key insight from this work is that a simple baseline (SimpleCIL) using only frozen PTM features with prototype classifiers can outperform state-of-the-art methods like L2P and DualPrompt by significant margins, demonstrating the inherent power of pre-trained representations for continual learning. The framework’s universality is evidenced by its compatibility with diverse architectures (ViTs and CNNs) and multiple PEFT approaches, while maintaining parameter efficiency and avoiding catastrophic forgetting through its adapt-once, merge-and-freeze strategy that balances the fundamental trade-off between generalizability and adaptivity in the pre-trained model era.
LAE (Learning-Accumulation-Ensemble) [36] proposes a unified continual learning framework that can transform any parameter-efficient tuning method into an effective continual learning approach. The framework positions prompting as merely one instantiation of PEFT and extends the concept to work with general parameter-efficient methods including Adapter, LoRA, and Prefix tuning. LAE operates through three key components: (1) Learning with calibrated speed to align the adaptation rates of different PET modules, addressing the issue that different PEFT methods vary in their knowledge acquisition speed which can lead to overfitting or insufficient plasticity; (2) Accumulation of multi-task knowledge where task-specific information learned by online PEFT modules is progressively accumulated into offline PET modules through sophisticated knowledge transfer mechanisms; and (3) Ensemble of two expert models constructed with online and offline PEFT modules that work collaboratively during inference to balance stability and plasticity. The framework’s key contribution lies in its ability to reshape any given PEFT method into a competitive memory-free continual learning approach while maintaining the computational efficiency advantages of parameter-efficient tuning. By providing theoretical foundations for why different PEFT methods require speed calibration and demonstrating superior performance across CIFAR100 and ImageNet-R benchmarks, LAE establishes a general paradigm that unifies the benefits of various parameter-efficient approaches under a single continual learning framework, making it broadly applicable across different model architectures and tuning strategies.